OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6.

jigglypuff314 (jigglypuff314):

first what is the slope of the line y = -x + 6?

OpenStudy (anonymous):

6?

jigglypuff314 (jigglypuff314):

close :) 6 would be the y-intercept y = mx + b where m=slope so what is the slope of y = -x+6

OpenStudy (anonymous):

-1 since there is an imaginary 1 before x?

jigglypuff314 (jigglypuff314):

correct :) then know that the slope of a line perpendicular to that one would be the flipped opposite of it for example the flipped opposite of -1/4 is 4

OpenStudy (anonymous):

so it would just be y=x+6?

jigglypuff314 (jigglypuff314):

not quite :) y = 1x + b temporarily plug in the point (-4, -2) so -2 = -4 + b solve for b

OpenStudy (anonymous):

b=-2

jigglypuff314 (jigglypuff314):

yes :) so for y = mx + b you now have m=1 and b=-2

OpenStudy (anonymous):

so now i solve tht?

jigglypuff314 (jigglypuff314):

mmm now you would just plug in :)

OpenStudy (anonymous):

-2=1x+(-2) right?

jigglypuff314 (jigglypuff314):

y=x -2

OpenStudy (anonymous):

oooo thanks :)